Phosphoric acid (H3,PO4) is a key ingredient in soft drinks that gives it a


sharper flavor. One way to produce this acid is with the aid of mineral


fluoroapatite (Ca5(PO4)3F) following this reaction:



Ca5(PO4)3F + 5H2SO4 → 3H3PO4 + 5CaSO4 + HF


If 24.5 g Ca5(PO4)3F is mixed with 89.7 g H2SO4



, how much H3PO4 will form?

1
Expert's answer
2021-11-11T08:24:02-0500

"Ca_5(PO_4)_3F + 5H_2SO_4 \u2192 3H_3PO_4 + 5CaSO_4 + HF"


number of moles of fluoroapatite = 24.5g/138.05g = 0.177 mol

number of moles of sulphuric acid = 89.7g/98g = 0.915 mol

number of reacting moles = 0.915/5 = 0.183 mol


the limiting reactant is fluoroapatite.

"\therefore" 0.183 moles of phosphoric acid is formed.


mass of the phosphoric acid = number of moles formed × number of stoichiometric moles × mass of phosphoric acid = 0.183 × 5 × 98 = 89.67g"."